OpenStudy (anonymous):

xtan27 = tan11sqrt(1600-x^2) can someone show steps to reaarage this to x^2+=?

OpenStudy (anonymous):

fan and medal

OpenStudy (anonymous):

xtan27 = tan11sqrt(1600-x^2) x^2tan^2(27) = (1600-x^2)tan^2(11) x^2tan^2(27) = 1600tan^2(11) - x^2tan^2(11) x^2[tan^2(27) + tan^2(11)] = 1600tan^2(11) so x^2 = 1600tan^2(11)/[tan^2(27) + tan^2(11)?

OpenStudy (alekos):

you need to work some of this out yourself. not get everyone else to do it

OpenStudy (anonymous):

i got it . man thnx alekos

OpenStudy (alekos):

good

ganeshie8 (ganeshie8):

Looks perfect!

OpenStudy (anonymous):

thnx

ganeshie8 (ganeshie8):

just take square root both sides and you're done!